OpenStudy (anonymous):

Evaluate the limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow+ \infty} = \frac{ 2x + 1 }{ \sqrt{x ^{2}+3} }\]

OpenStudy (anonymous):

This one needs a little manipulation to make it work.\[\lim_{x \rightarrow \infty}\frac{2x+1}{\sqrt{x^2+3}}=\lim_{x \rightarrow \infty}\frac{2x+1}{x \sqrt{1+\frac{3}{x^2}}}=\lim_{x \rightarrow \infty}\frac{2+\frac{1}{x}}{\sqrt{1+\frac{3}{x^2}}}\]You should be able to handle it from there.

OpenStudy (anonymous):

\[\lim_{x \rightarrow+ \infty} = \frac{ 2+ (+\infty) }{ \sqrt{1+(+\infty)} } = \infty \] right? @AnimalAin

OpenStudy (sidsiddhartha):

u need to know one thing that \[\lim_{x \rightarrow \infty}=\lim_{(1/x) \rightarrow 0}\]

OpenStudy (anonymous):

Not correct. Note that you are taking the limit as x approaches infinity. The reciprocal of a number going to infinity approaches zero. The answer is two, as noted earlier.

geerky42 (geerky42):

\[\lim_{x\rightarrow\infty} \dfrac{2x+1}{\sqrt{x^2+3}} = \lim_{x\rightarrow\infty}\dfrac{2x}{\sqrt{x^2}}\]Because when x approaches to infinity, terms with lower degree become less and less significant

OpenStudy (anonymous):

Ok now how it will compensate \[\infty \] in the equation?

OpenStudy (anonymous):

If you look at my work earlier, we divided out any value of x that wasn't in a denominator. As x goes to infinity, any bounded number divided by x (or x^2) will go to zero. After we get rid of those terms, we end up with 2/1 = 2.

OpenStudy (anonymous):

aha .. so the solution = 2 ?

OpenStudy (anonymous):

Can you give me a solution in a simplified way?

OpenStudy (anonymous):

I thought I did. You need to look carefully at the algebra I used to manipulate the problem in a way that made the values that go to infinity go away. Keep after it; it's not that hard. I mean, if I can do it, how hard could it be? Do math every day.

OpenStudy (anonymous):

In any case, thank you very much :*

OpenStudy (anonymous):

No sweat.

geerky42 (geerky42):

Well, I kind of did. For limit as x approaches to infinity, you just need to look and compare degree (highest power) of numerator and denominator. If degree in numerator is higher than denominator's, then limit is infinity. If degree in denominator is higher than numerator's, then limit is 0 If both have same degree, then limit is ratio of coefficient of leading term In this problem, both have same degree, so limit is 2/1 = 2

OpenStudy (anonymous):

Can we do that in the equation? \[\lim_{ء \rightarrow +\infty} = \frac{ 2x+1 }{ (x ^{2}+3)^{1/2} }\]

OpenStudy (campbell_st):

thats just the same question

geerky42 (geerky42):

Yes. You just focus on leading term in both numerator and denominator. you can see that \(\sqrt{x^2}\) = x, ratio of coefficient is 2/1 = 1. Does that make sense? Sorry if my English is bad lol

OpenStudy (anonymous):

so \[\lim_{x \rightarrow +\infty} \frac{ 2x + 1 }{ \sqrt{x ^{2} + 3} } = \lim_{x \rightarrow +\infty} \frac{ 2x+1 }{ \sqrt[x]{1+\frac{ 3 }{ x ^{2} }} }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow +\infty} \frac{ 2+\frac{ 1 }{ x } }{ \sqrt{1+\frac{ 3 }{ x ^{2}}} }\]

geerky42 (geerky42):

\[\Large \lim_{x\rightarrow\infty}\dfrac{2x+1}{\sqrt{x^2+3}} = \lim_{x\rightarrow\infty}\dfrac{2x}{\sqrt{x^2}} = \lim_{x\rightarrow\infty}\dfrac{2x}{x} = \lim_{x\rightarrow\infty} 2 = \boxed{2}\]

OpenStudy (campbell_st):

thats it.... now when you substitute \[\frac{1}{x} .. and..\frac{3}{x^2}\] both approach zero as x approaches infinity

geerky42 (geerky42):

Guess my approach isn't clear

OpenStudy (campbell_st):

so a simple method... was to factor the term inside the radical \[x^2 + 3 = x^2(1 + \frac{3}{x^2})\] then applying a radical law to the factored form you get \[x \sqrt{1 + \frac{3}{x^2}}\] now divide every term by x and you get \[\frac{2 + \frac{1}{x}}{\sqrt{1 + \frac{3}{x^2}}}\] so now in this form... as x approaches infinity... the value of the fractions approaches zero

OpenStudy (campbell_st):

which means you limit becomes \[\frac{2}{1}\]

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Thank you all :*